an object is released from a height h above the ground reaches the ground in 8sec the time taken by the object to cover the first half of the distance is?

Grade:12th pass

1 Answers

Naitik jain
6 years ago
u=0. Let a=g=10t=8 ut + 1/2at²=sS=1/2*10*8*8S=320 160= half distance travelled T for next half be x160=ux +1/2ax²160=0+1/2*10*x²X=under root 32Or 5.65seconds